The following data represent the muzzle velocity (in feet per second) of rounds fired from a 155-mm gun. For each round, twoArea in right tail Degrees of Freedom -Distribution Area in Right Tail 0.05 0.025 0.25 0.20 0.15 0.10 0.02 0.01 0.005 0.002519 20 0.001 0.860 1.000 1.064 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 0.000 0.687 0.686 0.686 0.685 0.685 0.(a) Is there a difference in the measurement of the muzzle velocity between device A and device B at the α = 0.01 level of​ significance? Note: A normal probability plot and boxplot of the data indicate that the differences are approximately normally distributed with no outliers. Let di = Ai − Bi.

(i) Identify the null and alternative hypotheses.

(ii) Determine the test statistic for this hypothesis test (t0 = ?). Round to two decimal places as needed.

(iii) Calculate avg(d) and sd. Round to six decimal places as needed.

(iv) Find the P-value (P = ?). Round to six decimal places as needed.

(b) Identify the lower and upper bounds for the 99% confidence interval. Round to six decimal places as needed.

(c) Draw a boxplot of the differenced data. Does this visual evidence support the results obtained in part​ (a)?
Answer #1
SAMPLE 1 SAMPLE 2 difference , Di =sample1-sample2 (Di - Dbar)²
790.4 791 -0.6 2.560
792.9 788.4 4.5 12.250
794.1 795.5 -1.4 5.760
792.9 788 4.9 15.210
794.9 797.6 -2.7 13.690
792.8 791.5 1.3 0.090
sample 1 sample 2 Di (Di - Dbar)²
sum = 4758 4752.00 6.000 49.560

Ho :   µd=   0                  
Ha :   µd ╪   0                  
                          
Level of Significance ,    α =    0.01                  
                          
sample size ,    n =    6                  
                          
mean of sample 1,    x̅1=   793.000                  
                          
mean of sample 2,    x̅2=   792.000                  
                          
mean of difference ,    D̅ =ΣDi / n =   1.0000                  
                          
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    3.148333
                          

std error , SE = Sd / √n =    3.1483   / √   6   =   1.2853      
                          
t-statistic = (D̅ - µd)/SE = (   1   -   0   ) /    1.2853   =   0.78
                          
Degree of freedom, DF=   n - 1 =    5                  
t-critical value , t* =    ±   4.0321   [excel function: =t.inv.2t(α,df) ]               
                          
p-value =        0.471746 [excel function: =t.dist.2t(t-stat,df) ]               
Decision:   p-value>α , Do not reject null hypothesis  

...............

sample size ,    n =    6          
Degree of freedom, DF=   n - 1 =    5   and α =    0.01  
t-critical value =    t α/2,df =    4.0321   [excel function: =t.inv.2t(α/2,df) ]      
                  
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    3.1483          
                  
std error , SE = Sd / √n =    3.1483   / √   6   =   1.2853
margin of error, E = t*SE =    4.0321   *   1.2853   =   5.1825
                  
mean of difference ,    D̅ =   1.000          
confidence interval is                   
Interval Lower Limit= D̅ - E =   1.000   -   5.1825   =   -4.182520
Interval Upper Limit= D̅ + E =   1.000   +   5.1825   =   6.182520

Chart Title 6 5 4 3 2 1 0 -1 -2 -3 -4 1


                  
      

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!